OpenStudy (anonymous):

x^3 -2 -2i sqrt(3) =0

OpenStudy (anonymous):

\[\begin{align*} x^3-2-2i\sqrt3&=0\\ x^3&=2+2i\sqrt3 \end{align*}\] Let's convert to polar so we can use DeMoivre's theorem: \[r^2=2^2+(2\sqrt3)^2=4+12=16~~\implies~~r=4\] \[\begin{cases}4\cos\theta=2\\4\sin\theta=2\sqrt3\end{cases}~~\implies~~\theta=\frac{\pi}{3}\] So you have \[\begin{align*}x^3&=2+2i\sqrt3\\ x^3&=4\left(\cos\frac{\pi}{3}+i\sin\frac{\pi}{3}\right)\\ x&=\sqrt[3]{4\left(\cos\frac{\pi}{3}+i\sin\frac{\pi}{3}\right)} \end{align*}\] Find all the third roots of the complex number.